# Water-rich sub-Neptunes and rocky super Earths around different Stars: Radii shaped by Volatile Partitioning, Formation, and Evolution

### v1

@@ 3.1 Evolutionary tracks
In order to understand the dependencies and behavior of the different model assumptions, it is illustrative to visualize the time evolution of three different planetary mass cases under the four different model assumptions. This covers qualitatively the different possibilities for the final states of initially water-rich planets. Here, we initialized the planets by using a total mass – luminosity relation of Linder et al. (2019), and H/He and water fractions typical for the formation output. The parameters are tabulated in Table 1. Furthermore, the planets are placed after at the same orbit with semi-major axis of 0.01 au around a 0.5 M⊙ star.
At time zero of the model, the Water Sequestration and Layered models distribute a fraction or all of the water contained in the envelope either to the core or it is layered on top of the core. The gaseous envelope is therefore depleted in heavy particles and adjusts itself to a larger radius. As soon as the simulation starts evolving in time, the larger radius leads to a higher photoevaporative mass loss rate. Thus, the planet looses mass and consequently shrinks in size more rapidly than the cases with a Mixed or Fractionation assumption.This can be observed for the intermediate and massive planet and the effect is more pronounced in the Layered scenario compared to the Water Sequestration scenario as all water instead of just a part, corresponding to the possible uptake of the core and magma ocean, is removed from the envelope.
Depending on the initial mass and structure assumption, the gaseous envelope can be lost completely. For the Layered model, the envelopes are removed in all cases. For the Water Sequestration model, the most massive planet retains its gaseous envelope while the intermediate and light planets transition to a secondary envelope stage. This can be seen from the yellow curve dipping to the size of a bare core before the numerical scheme establishes a secondary atmosphere due to the reservoir of water contained in the magma ocean being allowed to re-equilibrate with the atmosphere. As mentioned in the model description and will be discussed in Sect. 4.3, this approach is relatively crude and could be improved in the future. Generally, the outgassed atmosphere is almost entirely made of heavy elements and of relatively low-mass. Although it could be removed on timescale comparable to the lifetime of the system, the replenishment from the much larger magma ocean reservoir typically sustains it for planets with (Kite & Schaefer, 2021).
@@
The Fractionation model differs less than a percent in radius from the Mixed model. There are hints of weakly fractionating mass-loss at the end of the lifetime of the atmosphere on the light planet (300 kyr, dotted purple line) where the atmosphere is removed slightly faster in the Mixed model case. If an observation at that point would have been made, the case where fractionation is included would appear slightly larger. For the more massive planets, the effect of fractionation is to marginally increase over time which leads to percent-level smaller radii due to the slightly increased mean molecular weight. A numerical artifact is the initially increased radius for the Fractionation model output for the intermediate-mass planet. It originates due to differing timesteps just after initialization where mass loss rates are large and slight overshooting occurred for the Mixed model, revealing the level of precision obtainable with our approach. Generally, the model is not suited for radius comparisons below the 1% level.
@@ 3.2.1 Mixed model
We start by comparing the mass-radius relation of the synthetic planets using the nominal, mixed envelope assumption against the sample of observed planets in the top, left panel of Fig. 4. The synthetic distribution consists of volatile-rich planets with significant scatter in mass-radius relation distinct from a population of higher-density, rocky planets with a tightly correlated mass-radius relation. 333Not discussed in the rest of this study, we find a few planets below the main rocky composition relation, made of pure iron in the model, which are sometimes observed (Goffo et al., 2023). In our calculations, they tidally migrated towards the star and lost their silicate mantle due to photoevaporation, using a simple approximation of hydrodynamic or jeans escape of an atmosphere with an assumed mean molecular weight of 49 u.
@@ 3.2.2 Layered model
This leaves a large population of planets with pure water envelopes which is then protected from photoevaporative loss due to the higher atmospheric mean molecular weight compared to mixtures. Therefore, less planets populate the low-density parameter space at larger masses . Instead, many planets are located in the pure water world regime at around two R⊕.
@@ 3.2.4 Water Sequestration model
The results of the Water Sequestration model can be broadly placed between the Layered and the Fractionation model. A part of the water content separates from the gaseous envelope and is therefore not lost to photoevaporation. From Fig. 4, we see that down to a mass of 2 M⊕, large water reservoirs can remain on the planet. At the end of the evolution (see also Sect. 3.1), this water is present in the form of oxygen and hydrogen in the metallic core, in the magma ocean, as well as outgassed as steam atmosphere. Planets in this state are tightly correlated in mass-radius while planets with primordial H/He content are distributed to larger radii. The shape of the mass-radius distribution is similar to the output of the Layered model, however, the planets with steam atmospheres are significantly smaller than in that model. This is due to only a fraction of the initial water content remaining instead of the total amount as in the Layered model.
Compared to observations, the extension of water-rich planets to masses blow 3 M⊕ is not observed (see also Sect. 4.1). Furthermore, the model also predicts a population of H/He free planets with large water contents (i.e. water worlds) on a narrow mass-radius relation. While some observed planets populate this region of the parameter space, they are not as numerous as predicted by the model (see also the analysis of Parviainen et al. (2024) on observational evidence for water worlds). Furthermore, from Fig. 5, we find a population of these water wolds (with R⊕) at short orbital periods ( day) which is not matched by observations. The Mixed model reproduced this innermost distribution of planets better. However, the mass-radius distribution of larger sub-Neptunes is well-matched. In particular, there is less scatter than in the Mixed model and no planets with too low densities exist in the sub-Neptune regime. This indicates that lowering the water fraction in the gaseous envelope led to a better match with observations for planets keeping their primordial envelopes.
@@ 3.3 The radius valley as a function of stellar mass
Here, we investigate the dependency of the emerging radius valley on stellar mass. In Fig. 6, we show for the Mixed model the full synthetic data without an observational bias applied for different discrete stellar masses. Since the previously introduced bias is applicable to planets characterized in mass and radius, we opt to not use an estimate of the bias in this radius comparison (irrespective of planetary masses) to investigate the actual trends predicted by the model. Complementarily, we will discuss the radius valley trends using the populations around Solar mass stars with applied Kepler observational bias in Sect. 3.4. To support the discussion here, we also show a summary of fits done for observed or synthetic data in Fig. 7 and mass-radius relations resulting from the Mixed model for the different stellar mass bins in Fig. 8. The results of the other models are presented in the same fashion in Appendix C. In each panel of Figs. 6, 16, and 17, gapfit (Loyd et al., 2020) was used with bootstrapping to infer a gap in the region of the observed gap, that is, using planets from 1 day to 200 days orbital period for stellar masses larger than 0.1 M⊙ and from 0.6 to 20 days for the 0.1 M⊙ case. The gap is fit using a power-law
@@
The search for a gap is performed within 0.15 dex of , with an initial slope guess of , and a Kernel width of 0.15.
@@
The synthetic data at low stellar mass for the Mixed model shows a similar gap location at in agreement with Luque & Pallé (2022). We can see a trend with stellar mass: the locus of the valley at orbital period shifts upward for late M dwarfs before converging to for stellar masses larger than . This kind of threshold stellar mass is similar to the mass at which the small planet population no longer increases significantly in mass, as discussed in Burn et al. (2021). In that work, we attributed this to the fact that growth of small planets is no longer limited by the amount of solids in the disk at stellar masses larger than which we might recover here since growth determines the initial mass and thus indirectly the radius of the Super-Earth population.
@@
The observed radius valley indicates a dependency on stellar mass around FGK-type stars based on Kepler and K2 data (Berger et al., 2020; Petigura et al., 2022; Ho & Van Eylen, 2023). The reported logarithmic slopes of the radius valley location with stellar mass range from 0.1 to 0.4 with likely values around 0.2. However, for M stars, Luque & Pallé (2022) report a slope consistent with no stellar mass dependency . Although it is likely not the relevant functional form, we fit a power-law to the inferred radius valley locations for the five presented stellar mass bins at 1 day orbital period and obtain , which agrees with the observed trend for M dwarfs (Luque & Pallé, 2022) but is flatter than found for FGK stars where values around 0.2 are reported (Petigura et al., 2022; Ho & Van Eylen, 2023). From Fig. 7, we see that also the data by Petigura et al. (2022) shows a flattening of for the two higher stellar mass bins.
@@
While the locus of the valley increases, our inferred slopes with orbital period hints at a flatter dependency with increasing stellar mass. The values range from at to -0.07 for Solar mass stars. This is comparable to other theoretical works relying on photoevaporation predictions from hydrodynamic models instead of simple energy- or recombination-limited escape models (the latter resulting in steeper slopes Mordasini, 2020; Affolter et al., 2023). Although the error from the bootstrap method is small, we expect that the flattening is not based on enough statistical data from the model to confirm it as a significant trend. The flattening could be influenced from below, by growing more massive super-Earths at larger orbital periods around more massive stars, or from the top by less efficient photoevaporation and thus smaller radii of volatile-rich planets at fixed orbital period around lower-mass stars. Our results remain for all stellar masses in the regime of negative slopes indicating that photoevaporation is the dominant effect shaping the valley. The compilation of observational data (Fig. 7) does not show a significant trend with the exception of the inferred slope by Luque & Pallé (2022). However, the theoretical work of Venturini et al. (2024) predicts a steeper slope around more massive stars which was not recovered here.
@@
Ho et al. (2024) and Bonfanti et al. (2024) gave indications for a more populated, radius valley around low-mass stars. Similarly, Cloutier & Menou (2020) already noted a narrowing of the radius valley towards lower stellar masses. Apart from the 0.1 M⊙ case where the statistics becomes poor, we find a similar trend in Fig. 6. This trend was also recovered and discussed in Venturini et al. (2024). From the detailed discussion of mass-radius relations (Sect. 4.1), we interpret part of it being linked to the low-mass tail of the volatile distribution extending into the radius valley (but not the density valley, see Luque & Pallé, 2022; Venturini et al., 2024). We note that the M star sample probes the planetary population at lower irradiations where condensation could occur, but we excluded this region when we plot biased populations. We leave a more detailed analysis of the pollution of the valley to future works.
@@
In addition to the data, we show an estimate of the runaway greenhouse limit (Sect. 2.3) as gray line in Fig. 6. The fact that for low-mass stars, this limit is not necessarily outside the regime of interest further motivates the investigation of a model which does not assume that water and H/He mix, that is, the Layered model. Thus, in Fig. 16, we show the resulting unbiased planetary population with the previously introduced compositional color code using the Layered model. The most striking effect is that the population of H/He-free sub-Neptunes is more numerous and extends to longer orbital periods than in the Mixed model. If the assumption that no water enters the low-density envelope is made, also planets at larger orbital periods can lose their H/He content. This also leads to a more abrupt transition from H/He free to H/He-rich planets at a threshold radius for a given orbital period. Therefore, a second radius valley emerges between H/He-rich and H/He-free planets at . So far and to our knowledge, no reports on such a valley have been published. However, a visual indication of such an under-density can be made out in the data of Ho & Van Eylen (2023) for Solar mass stars, their Fig. 1, although it does not seem to be of high enough significance. We further note that in the Solar mass data, the observable region is well within the runaway greenhouse limit where we expect mixing.
@@
The algorithm could determine a radius valley for all but the 0.1 M⊙ case for the Layered model. It lies between rocky planets and water worlds and at too low radii compared to observations (Fig. 7) because pure water envelopes instead of mixed-composition ones are less prone to being removed by photoevaporation (see e.g. Mordasini, 2020, for a parameter study). However, the distance to the observed valley is comparable to the Mixed model. This seems to naively suggest that reality lies between the modeled assumptions.
For the Water Sequestration model, no radius valley can be fitted conclusively around the three lowest stellar mass bins. Although there is a transition in density space, the existence of steam worlds with low mass (M⊕) makes this density valley undetectable in radius space. This can also be seen when comparing the mass-radius relations shown in Fig. 8 with Fig. 15.
@@ 3.4 The radius valley as seen by Kepler
The ability of the different models to reproduce the observed radius valley can also be tested against the Kepler dataset. As mentioned in Sect. 2.3, we constructed a set of synthetic systems as the Kepler spacecraft would observe it. Given by the sensitivity of Kepler, we restrict the selection of simulations to those around Solar mass stars. In Fig. 9, the resulting cumulative distributions of the synthetic planetary radii can be compared to the observed ones. We further added the distribution by Burn et al. (2024) since several assumptions on the initial conditions were changed although the Mixed model is physically similar to the model used in that work.
We find a reduced number of sub-Neptunes in the Mixed model compared to Burn et al. (2024) in disagreement with observations. The high-radius tail of sub-Neptunes, that is, the radius cliff, is best matched in the Mixed model. The radius valley location, also discussed in Sect. 3.3 where a radius valley location at 2.4 R⊕ was found, is shifted to too large radii in the Mixed model. This implies that the sub-Neptunes in the Mixed model loose their envelopes too rapidly. However, we caution that the sub-Neptune to rocky planet ratio is sensitive to uncertain formation model assumptions (production of rocky planetesimals and planets), as well as to the application of the observational bias. The locus of the radius valley is less sensitive to these considerations.
From Fig. 9, we find that Layered model reproduces better the observed radius valley location, although shifted to too low radii for lower stellar masses (see Fig. 7). Furthermore, it has a more favorable sub-Neptune to rocky planet ratio compared to the Mixed model. From this comparison on planetary radii only, one might conclude that it should be favored over the Mixed model. In contrast, the Water Sequestration model lacks the distinct radius valley feature.

### v2

@@ 3.1 Evolutionary tracks
In order to understand the dependencies and behavior of the different model assumptions, it is illustrative to visualize the time evolution of three different planetary mass cases under the four different model assumptions. This covers qualitatively the different possibilities for the final states of initially water-rich planets. Here, we initialized the planets by using a total mass – luminosity relation of Linder2019, and H/He and water fractions typical for the formation output. The parameters are tabulated in Table 1. Furthermore, the planets are placed after at the same orbit with semi-major axis of 0.01 au around a 0.5 M⊙ star.
At time zero of the model, the Water Sequestration and Layered models distribute a fraction or all of the water contained in the envelope either to the core or it is layered on top of the core. The gaseous envelope is therefore depleted in heavy particles and adjusts itself to a larger radius. As soon as the simulation starts evolving in time, the larger radius leads to a higher photoevaporative mass loss rate. Thus, the planet looses mass and consequently shrinks in size more rapidly than the cases with a Mixed or Fractionation assumption. This can be observed for the intermediate and massive planet and the effect is slightly more pronounced in the Layered compared to the Water Sequestration scenario as all water, instead of just a part, corresponding to the possible uptake of the core and magma ocean, is removed from the envelope.
Depending on the initial mass and structure assumption, the gaseous envelope can be lost completely. For the Layered model, the envelopes are removed in the intermediate and low mass cases and replaced by pure water envelopes from the layer below. The intermediate-mass case goes through an inflation phase shortly before the H/He mass is completely lost just after . This is not completely physically correct as the AQUA layer is assumed to remain fully adiabatic in our simplified treatment during the evolution phase. In reality, it should become partially radiative. The outcome is nevertheless as expected such that the H/He is fully lost and a new consistent hydrostatic envelope structure (incl. a radiative layer) made of pure water is initialized.
Similarly, for the Water Sequestration model, the most massive planet retains its gaseous envelope while the intermediate and light planets transition to a secondary envelope stage via the equilibration of the interior with the atmosphere, that is, due to gradual outgassing of water. It leads to an increase in as can be seen in the bottom plot. As mentioned in the model description and will be discussed in Sect. 4.3, this approach is relatively crude and could be improved in the future. For lower mass planets (here the intermediate and low-mass case), the resulting atmosphere is almost entirely made of heavy elements . Although it could be removed on timescale comparable to the lifetime of the system, the replenishment from the much larger magma ocean reservoir typically sustains it for planets with (Kite2021). This result also already hints at a bifurcation of with either low values, where hydrogen and oxygen are still stored in the core, or high for planets which underwent efficient mass loss and significant outgassing.
@@
The Fractionation model differs less than a percent in radius from the Mixed model. There are hints of weakly fractionating mass-loss at the end of the lifetime of the atmosphere on the light planet (300 kyr, dotted purple line) where the atmosphere is removed slightly faster in the Mixed model case. For the more massive planets, the effect of fractionation is to marginally increase over time which leads to percent-level smaller radii due to the slightly increased mean molecular weight.
@@ 3.2.1 Mixed model
We start by comparing the mass-radius relation of the synthetic planets using the nominal, mixed envelope assumption against the sample of observed planets in the top, left panel of Fig. 4. The synthetic distribution consists of volatile-rich planets with significant scatter in mass-radius relation distinct from a population of higher-density, rocky planets with a tightly correlated mass-radius relation.
@@ 3.2.2 Layered model
Since the H/He loss is efficient, this leaves a large population of planets with pure water envelopes which is then protected from photoevaporative loss due to the small radii caused by the higher atmospheric mean molecular weight compared to mixtures. Therefore, less planets populate the low-density parameter space at larger masses . Instead, many planets are located in the pure water world regime at around two R⊕.
@@ 3.2.4 Water Sequestration model
The results of the Water Sequestration model can be broadly placed between the Layered and the Fractionation model. A part of the water content, usually the majority by mass, separates from the gaseous envelope and is therefore not lost to photoevaporation. From Fig. 4, we see that down to a mass of 1 M⊕, large water reservoirs can remain on observable planets. At the end of the evolution (see also Sect. 3.1), this water is present in the form of oxygen and hydrogen in the metallic core, in the magma ocean, as well as outgassed as steam atmosphere. Planets in this state are tightly correlated in mass-radius while planets with primordial H/He content are distributed to larger radii. The shape of the mass-radius distribution is similar to the output of the Layered model, however, the planets with steam atmospheres are significantly smaller than in that model. This is due to only a fraction of the initial water content being present in gaseous form and reduced total water mass compared to the Layered model.
Compared to observations, the extension of water-rich planets to masses below 3 M⊕ is not observed (see also Sect. 4.1). Furthermore, from Fig. 5, we find a qualitatively worse match to observations with a too narrow radius valley at short orbital periods for Sequestration model compared to the Mixed one. We refrain from further interpretation of the differences because the observed and modeled distributions do not match well in period space in general which is likely a shortcoming of the formation model and not necessarily related to the long-term evolution.
Despite these differences, the mass-radius distribution of planets with masses above 3 M⊕, including larger sub-Neptunes, is well-matched in the Water Sequestration model. In particular, no planets with too low densities exist in the sub-Neptune regime while such planets existed in the Mixed model output. The outcome of the model differs in this regime because the same planet (similar to the high-mass case shown in Fig. 3) initially has a lower . The even lower density planets than the Mixed case then undergo significant mass loss before outgassing of metals enhances their . Combined, this leads to higher density planets. Furthermore, the initial is more uniform across planets as it is regulated by the core and mantle reservoir.
Lastly, no massive bare rocky planets are predicted in agreement with observations. This indicates that lowering the water fraction in the gaseous envelope led to a better match compared to observations for planets with M⊕. This is also the regime where the sequestration scenario is most appropriate due to higher accretion energies and thus hotter cores which are more likely to be completely molten.
@@ 3.3 The radius valley as a function of stellar mass
Here, we investigate the dependency of the emerging radius valley on stellar mass. In Fig. 6, we show for the Mixed model the full synthetic data without an observational bias applied for different discrete stellar masses. Since the previously introduced bias is applicable to planets characterized in mass and radius, we opt to not use an estimate of the bias in this radius comparison (irrespective of planetary masses) to investigate the actual trends predicted by the model. Complementarily, we will discuss the radius valley trends using the populations around Solar mass stars with applied Kepler observational bias in Sect. 3.4. To support the discussion here, we also show a summary of fits done for observed or synthetic data in Fig. 7 and mass-radius relations resulting from the Mixed model for the different stellar mass bins in Fig. 8. The results of the other models are presented in the same fashion in Appendix C. In each panel of Figs. 6, 16, and 17, gapfit (Loyd2020) was used with bootstrapping to infer a gap in the region of the observed gap, that is, using planets from 1 day to 200 days orbital period for stellar masses larger than 0.1 M⊙ and from 0.6 to 20 days for the 0.1 M⊙ case. The gap is fit using a power-law
@@
The search for a gap is performed within 0.15 dex of , with an initial slope guess of , and a Kernel width of 0.15 dex.
@@
The synthetic data at low stellar mass for the Mixed model shows a similar gap location at in agreement with Luque2022. We can see a trend with stellar mass: the locus of the valley at orbital period shifts upward for late M dwarfs before converging to for stellar masses larger than . This conclusion also holds when analyzing the populations as a function of integrated, received EUV or X-ray flux instead of orbital period. This kind of threshold stellar mass is similar to the mass at which the small planet population no longer increases significantly in mass, as discussed in Burn2021. In that work, we attributed this to the fact that growth of small planets is no longer limited by the amount of solids in the disk at stellar masses larger than which we might recover here since growth determines the initial mass and thus indirectly the radius of the Super-Earth population.
@@
The observed radius valley indicates a dependency on stellar mass around FGK-type stars based on Kepler and K2 data (Berger2020a; Petigura2022; Ho2023). The reported logarithmic slopes of the radius valley location with stellar mass range from 0.1 to 0.4 with likely values around 0.2. However, for M stars, Luque2022 report a slope consistent with no stellar mass dependency . Although it is likely not the relevant functional form, we fit a power-law to the inferred radius valley locations for the five presented stellar mass bins at 1 day orbital period and obtain , which agrees with the observed trend for M dwarfs (Luque2022) but is flatter than found for FGK stars where values around 0.2 are reported (Petigura2022; Ho2023). From Fig. 7, we see that also the data by Petigura2022 shows a flattening of for the two higher stellar mass bins.
@@
While the locus of the valley increases, our inferred slopes with orbital period hints at a flatter dependency with increasing stellar mass. The values range from at to -0.07 for Solar mass stars. This is comparable to other theoretical works relying on photoevaporation predictions from hydrodynamic models instead of simple energy- or recombination-limited escape models (the latter resulting in steeper slopes Mordasini2020; Affolter2023). Although the error from the bootstrap method is small, we expect that the flattening is not based on enough statistical data from the model to confirm it as a significant trend. The flattening could be influenced from below, by growing more massive super-Earths at larger orbital periods around more massive stars, or from the top by less efficient photoevaporation and thus smaller radii of volatile-rich planets at fixed orbital period around lower-mass stars. Our results remain for all stellar masses in the regime of negative slopes indicating that photoevaporation is the dominant effect shaping the valley. The compilation of observational (Fig. 7) does not show a significant trend. However, the theoretical work of Venturini2024 predicts a steeper slope around more massive stars which was not recovered here.
@@
Ho2024 and Bonfanti2024 gave indications for a more populated, radius valley around low-mass stars. Similarly, Cloutier2020 already noted a narrowing of the radius valley towards lower stellar masses. Apart from the 0.1 M⊙ case where the statistics becomes poor, we find a similar trend in Fig. 6. This trend was also recovered and discussed in Venturini2024. From the detailed discussion of mass-radius relations (Sect. 4.1), we interpret part of it being linked to the low-mass tail of the volatile distribution extending into the radius valley (but not the density valley, see Luque2022; Venturini2024). We note that the M star sample probes the planetary population at lower irradiations where condensation could occur, but we excluded this region when we plot biased populations. We leave a more detailed analysis of the pollution of the valley to future works.
@@
In addition to the data, we show an estimate of the runaway greenhouse limit (Sect. 2.3) as gray line in Fig. 6. The fact that for low-mass stars, this limit is not necessarily outside the regime of interest further motivates the investigation of a model which does not assume that water and H/He mix, that is, the Layered model. Thus, in Fig. 16, we show the resulting unbiased planetary population with the previously introduced compositional color code using the Layered model. The most striking effect is that the population of H/He-free sub-Neptunes is more numerous and extends to longer orbital periods than in the Mixed model. If the assumption that no water enters the low-density envelope is made, also planets at larger orbital periods can lose their H/He content. This also leads to a more abrupt transition from H/He free to H/He-rich planets at a threshold radius for a given orbital period. Therefore, a second radius valley emerges between H/He-rich and H/He-free planets at . So far and to our knowledge, no reports on such a valley have been published. However, a visual indication of such an under-density can be made out in the data of Ho2023 for Solar mass stars, their Fig. 1, although it does not seem to be of high enough significance. We further note that in the Solar mass data, the observable region is well within the runaway greenhouse limit where we expect mixing.
@@
The algorithm could determine a radius valley for all but the 0.1 M⊙ case for the Layered model. It lies between rocky planets and water worlds and at too low radii compared to observations (Fig. 7) because pure water envelopes instead of mixed-composition ones are less prone to being removed by photoevaporation (see e.g. Mordasini2020, for a parameter study). However, the distance to the observed valley is comparable to the Mixed model. This seems to naively suggest that reality lies between the modeled assumptions.
For the Water Sequestration model, the algorithm also successfully fitted a radius valley for the stellar masses above 0.3 M⊙. In contrast to the previous model results, the core-regulated results do not depend on stellar mass with equal resulting for the different stellar masses. Furthermore, is flatter than in the models driven by photoevaporation. Visual inspection of the results in Fig. 17 indicates a more populated radius valley related to the water-rich sub-Neptune branch extending to lower planetary masses and radii (see Fig. 4). We also caution that the radius valley in this model emerges because planets with water in their interior efficiently outgas and instantaneously establish an atmosphere in equilibrium with the interior. In a more realistic scenario, there would be a timescale associated to outgassing which would not at all orbital periods be smaller than that of the evaporative loss timescale, thus the radius valley location, which still separates in this model the completely dry from water-rich planets with small outgassed atmospheres, would change or might not emerge at all.
For the Water Sequestration model, no radius valley can be fitted conclusively around the two lowest stellar mass bins. Although there is a transition in density space, the existence of steam worlds with low mass (M⊕) makes this density valley undetectable in radius space. This can also be seen when comparing the mass-radius relations shown in Fig. 8 with Fig. 15 and again supports the aforementioned interpretations of the radius valley in Venturini2024 and Luque2022 although here with a stronger overlap due to the low-mass volatile-rich planets.
@@ 3.4 The radius valley as seen by Kepler
The ability of the different models to reproduce the observed radius valley can also be tested against the Kepler dataset. As mentioned in Sect. 2.3, we constructed a set of synthetic systems as the Kepler spacecraft would observe it. Given by the sensitivity of Kepler, we restrict the selection of simulations to those around Solar mass stars. In Fig. 9, the resulting cumulative distributions of the synthetic planetary radii can be compared to the observed ones. We further added the distribution by Burn2024 since several assumptions on the initial conditions were changed although the Mixed model is physically similar to the model used in that work.
We find a reduced number of sub-Neptunes in the Mixed model compared to Burn2024 in disagreement with observations. The high-radius tail of sub-Neptunes, that is, the radius cliff, is best matched in the Mixed model. The radius valley location, also discussed in Sect. 3.3 where a radius valley location at 2.4 R⊕ was found, is shifted to too large radii in the Mixed model. This implies that the sub-Neptunes in the Mixed model loose their envelopes too rapidly. However, we caution that the sub-Neptune to rocky planet ratio is sensitive to uncertain formation model assumptions (production of rocky planetesimals and planets, see also discussion in Sect. 4.4.1), as well as to the application of the observational bias. The locus of the radius valley is less sensitive to these considerations.
From Fig. 9, we find that the Layered model reproduces better the observed radius valley location, although shifted to too low radii for lower stellar masses (see Fig. 7). Furthermore, it has a more favorable sub-Neptune to rocky planet ratio compared to the Mixed model. From this comparison on planetary radii only, one might conclude that it should be favored over the Mixed model. In contrast, the Water Sequestration model lacks the distinct radius valley feature in the cumulative distribution.